# Coherently Coupled Carrier and Phonon Dynamics in Elemental Tellurium Probed by Few-Femtosecond Core-Level Transient Absorption

### v1

@@ III.3 Temporal Dynamics of the A1g Phonon and Carriers
We now turn to the temporal dynamics of the three SVD components that track the kinetic processes following the excitation. In Fig. 4 the temporal vectors exhibit distinct dynamics that encode the edge shift, broadening, and carriers in detail. Focusing on the exponential rises and decays, the temporal dependence of the edge shift contribution (labeled phonon) has a sharp initial rise within the first few femtoseconds of the optical excitation followed by a slower onset over a picosecond. The fast rise can be ascribed to a rapid structural rearrangement driven by the depopulation of lone pair states near the Fermi level to anti-bonding states in the conduction band [8]. The rapid destabilization of the ground state potential energy surface results in a displacive force along the A1g phonon coordinate that decreases the helical chain radius by shifting the equilibrium position (). This increases the symmetry of the Te structure, inducing a displacive excitation of a coherent phonon (DECP) [15]. The change in local environment is what gives rise to an absorption edge shift due to changes in the 4d binding energy and decrease in the band-gap. The picosecond rise is typical of an incoherent growth in the acoustic phonon population that emerges from decay of the optical phonons [32]. The decay of the acoustic phonons is heavily suppressed by the very low lattice thermal conductivity of Te [7] and occurs over the course of milliseconds, as determined from the presence of a large heat background in our TA spectra. The decay we measure is thus due to the relaxation of the excited state potential energy surface over tens of picoseconds.
@@
Notably, we observe oscillations in both the phonon and hot carrier signals. While observation of the A1g phonon in Te is well established, oscillations in the hot carrier signal at the same frequency is similar to the back-coupling of the electronic temperature to the lattice that has been observed in semi-metallic Bi [18], Sb [19], and CDW system TaSe2 [20]. The hot carrier signal follows the A1g phonon motion, indicating the coupling arises from movement along the A1g phonon coordinate, corresponding to changes in the Te chain radius. The phase shift implies that when the Te helices are symmetrized (phonon maximum excursion), the carriers have a local temperature minimum. We discuss the potential source of this phase shift and quantify the temporal dynamics in the following section IV.
@@ IV.1 Multi-Temperature Model
To quantify the dynamics of the phonons and carriers the three SVD traces were globally fit using (Equation 2) which corresponds to a multi-temperature model implemented using the Minuit2 library [34]. A Gaussian instrument response function, , is retrieved by an external measurement and fixed in the fitting routine. In this model, the hot carriers are an electronic temperature () bath that may transfer energy to the lattice via an electron-phonon scattering rate (). The electronic temperature receives energy via the initially non-thermal carriers following photo-excitation through fast and slow scattering channels consistent with the biexponential rise of the hot carriers (e = electronic and p = phonon). The phonon trace is described by both the tellurium atom displacement from the equilibrium position along the A1g coordinate () and the acoustic phonon temperature (). The acoustic phonon temperature is fixed to rise at twice the rate of the coherent phonon decay () as the coherent optical phonon has been shown to decay via a Klemen’s like mechanism [32] through an anharmonic coupling of the lattice displacement to degenerate acoustic modes [35]. The rate for the initial displacement of the tellurium atoms is described by and relaxation back to the ground state potential energy surface (). Finally, the coherent energy exchange between the coherent optical phonon and electronic temperature is described following the approach of Drescher et al. [19] with the A1g frequency (), coherent phonon phase (), and a frequency chirp of (). The transient broadening , while not shown in Equation 2, is fit for completeness and follows the acoustic phonon temperature and phenomenologically the electronic component of the coherent energy exchange.
@@
At a carrier density of 1.5, the A1g phonon frequency retrieved is = 3.16(1) THz compared to the equilibrium 3.6 THz measured by Raman spectroscopy (see black curves in Fig. 4). The softening of the A1g mode is not surprising as the frequency is known to be dependent upon both carrier density [16, 9], and temperature in the form of anharmonic phonon-phonon coupling [9]. However, when subjected to a few percent excitation of the valence electrons, an increase in temperature is expected to signifigantly blue shift the A1g phonon frequency [9]. The 12.2 % decrease in frequency of the A1g motion measured with the XUV can then be attributed to a decrease in bond strength as the non-bonding lone pairs at the Fermi level are nearly instantaneously excited into the conduction band anti-bonding orbitals. This is in good agreement with the measured frequency determined from optical reflectivity pump-probe spectroscopy of polycrystalline Te [16, 9]. Similarly, at a carrier density of 2 the A1g phonon frequency is found to be 3.6(3) THz, closer to the equilibrium value (See Appendix E) with no apparent decay. The coherent phonon lifetime of ps and chirp of 1.02(7) ps-2 at 1.5 is typical of high fluence [9], while the coherent phonon phase is determined to be -0.03(1) rad and is discussed further below in Section IV.2.
@@
The rise () and decay () of the Te displacement from the equilibrium position () is now considered. The lifetime of the decay of from the excited state potential energy surface is ps and is much slower than the hot carrier decay. This lifetime of the Te displacement in the excited state potential energy surface is nearly identical to the carrier Auger recombination lifetime (31 ps) at the band-edge measured via optical pump-probe spectroscopy in 100 nm thick flakes [14] and the photo-induced degeneracy lifting of the conduction band minimum (H6 band) relaxation time (30 ps) [13]. In the original formulation of the displacive excitation of a coherent phonon model it is unclear whether the displacement is dependent upon the carrier population, , or if only the electronic temperature contributes to the displacement [15]. It is notable that in the multi-temperature model follows a timescale comparable to Auger recombination of fully relaxed carriers, such that . This could potentially explain why the displacement of the Te equilibrium position is much larger than in A7 semimetals [19, 18] where it is proposed to follow the electronic temperature. Additionally, when measured at a carrier density of , has no measurable decay out to the maximum time delays tested ( 2 ps). This is consistent with lengthening of the carrier lifetime as Auger recombination is reduced at low carrier densities, and multi-phonon mediated recombination is only expected to occur over 200 ps [14]. Moreover, there is a fast rise () to , which while nearly instantaneous with the sub 5 fs pump excitation, is found to be fs in a measurement with 660 as steps (see Appendix F). The delayed response of the Te equilibrium position displacement could be due to a lifetime dependence of the displacive force and non-equal contributions of the carriers to the excited state potential energy surface [36].
@@
The dynamics of the hot carriers are examined in Fig. 4. Upon initial excitation, the non-thermal carriers far from-equilibrium will quickly redistribute their energy through various scattering channels as they thermalize. The biexponential rise of the multi-temperature model for the hot carriers includes a fast term on the order of the 6.4(5) fs instrument response, and a slower 0.15(1) ps rise. Both timescales are much faster than previously measured inter-valley scattering that necessitates a population of high momentum acoustic phonons [37, 13]. Thus the primary source is likely intra-valley scattering towards the Fermi level in the form of energy-relaxation that is typically on the order of 100 fs [38, 39]. It is however possible that both carrier-carrier and carrier-optical phonon scattering are the dominant mechanisms for thermalization [40, 26]. While carrier-carrier scattering can only redistribute energy within the electronic system, the maximum energy exchange is on the order of the 800 nm photon energy. This will lead to much faster thermalization than through optical phonon emission that is constrained to the tens of meV transfer per scattering event. The fast rise () is thus attributable to carrier-carrier scattering, while the slower thermalization channel can be attributed to electron-phonon coupling, which is similar to the 0.3 ps intra sub-band cooling time measured with photoelectron spectroscopy [11]. The discrepancy could be due to the fact that in the analysis the hot carrier signal included both holes and electrons and that there are differences in scattering rates for holes versus electrons. The hot carrier decay () has a relaxation time of 1.60(2) ps that is similar to the timescale (1.3 ps [37], 1.7 ps [11]) for the transfer of high-energy carriers to the lowest energy sub-bands of the H-valley. In this process, for the hot carrier decay, the electronic temperature is reduced as optical-phonon emission through electron-phonon coupling that transfers energy to the lattice. The hot carriers determined from the SVD do not include the cooled lowest energy carriers that share the long lifetime of the Te equilibrium position displacement. Similarly, the non-thermal carriers are shown in the residual TA spectra alongside the cooled carriers (See Fig. 9 of Appendix D).
IV.2 Coupled Carrier and A1g Phonon Dynamics
@@
We now consider the coherent modulations of that are observed. Following the absorption of an ultrashort optical laser pulse by Te, there is a rapid increase in the total energy of the electronic degrees of freedom. This abrupt increase in the carrier population () and electronic temperature () contributes to excite the A1g phonon mode and is well explained by the DECP model [15] and additional theory by O’Mahony et al. [36] While Te is a semiconductor in the ground state, the displacement along the A1g phonon coordinate leads to a symmetric Dirac metal through a light-induced phase transition [8]. This phase transition can similarly be achieved by depopulating the doubly degenerate VBM lone-pair orbitals with shear strain to the CBM [3]. The increase in metallic character when Te becomes more symmetric by reducing the helical chain radius should lead to an increase in the electronic heat capacity [41] as the density of states is increased at the Fermi level [42]. If we consider the temporal evolution of the hot carriers in Fig. 4, the phase offset is close to with respect to the phonon excursion. When the phonon displacement is maximum and Te most metallic, the electronic heat capacity is maximized, leading to a lower temperature of the hot carriers. This similarly follows the thermodynamic model of Giret et al. [17] that describes a coupling between the electronic temperature and lattice motion in Bi as an isentropic conservation of entropy. The same model has been applied to interpret XUV pump-probe results in Sb [19] and Bi [18]. While the thermodynamic model anticipates that the mass of the atoms scales the change in electronic temperature, the temperature oscillations in Sb [19] with similar mass to Te, are smaller. This implies the oscillations observed in electronic temperature of Te are more nuanced possibly due to the greater equilibrium position displacement. In the vicinity of the semiconductor-to-metal phase transition [8], the dependence of the electronic heat capacity on could exhibit a large variation. This is anticipated as there are numerous band crossings at the Fermi level that emerge upon symmetrizing Te [8, 42]. While an indirect band gap is closed at the A-valley for a 14 pm displacement [42], complete symmetrizing of Te at 26 pm additionally closes the gap at the H-valley [8]. Under the conditions anticipated at our fluence, an opening and closing (maximum phonon displacement) of the gap will heavily influence the electronic heat capacity by transiently changing the number of available states in proximity to the Fermi level. A large excursion radius of the A1g phonon in Te could amplify this effect. The same dependence on could similarly emerge in the form of a dynamical coupling between the carriers and transiently changing band structure [42] in which the carriers rapidly redistribute, rather than remain in their bands. In addition to the change in electronic temperature, this non-adiabatic change in band occupations would produce oscillations in the hot carriers at the same frequency as the A1g phonon motion. While we cannot discern the source from our experiment, calculations of the electronic heat capacity changes from the light-induced semiconductor to metal phase transition may prove useful in the future in understanding the observations.
@@
Finally, we consider a Fourier analysis of the phonon-induced oscillations shown in Fig. 5. This is accomplished by subtracting the exponential contributions to the TA captured by the SVD and adding back the residual TA of Fig. 9. Further, a decomposition of the spin-orbit split 4d core-levels is done following the methodology of Zuerch et al. [29] to eliminate spectral congestion between the and to conduction band transitions. Fig. 5(a), the power spectrum () with only the phonon-induced oscillations shows a maximum intensity near the Fermi level of 40.4 eV, while above 42 eV, weaker features are associated with the appearance of phonon-induced features from the large lattice distortion found in Fig. 3(b). These spectral features that appear far from the band-edge due to a large could be a useful spectral fingerprint of the Te electronic structure to compare with theoretical calculations of the XUV absorption for the semiconducting and metallic phases. The peak intensity is found around the A1g mode frequency, however, there is considerable asymmetry and breadth in the spectrum with a shoulder at 2.3 to 2.5 THz. While the chirp of the phonon should shift the asymmetry towards higher frequencies, the appearance of a feature towards lower frequencies is consistent with the excitation of the much weaker E1TO phonon mode whose equilibrium frequency is 2.7 THz [13]. This was observed only at low temperatures previously in Te using an optical-pump probe method at 2.45 THz [9]. It is however possible the XUV core-level probe is more sensitive to the E1TO mode depending on the core-excited potential energy surface; morever, the shorter laser pulse could increase the amount of coherence, and our higher excitation fluence could increase the amount of excitation of the E1TO phonon mode at room temperature. Similarly, an asymmetric E1g mode has been observed in MoTe2 where it is only a fraction of the A1g intensity [26]. We thus consider both the A1g and E1TO modes in a phase analysis of the coherent phonon motions.
@@
In Fig. 5(b), the spectral phase is shown for the A1g mode in vicinity of the Fermi level. In agreement with the phase of rad retrieved from the multi-temperature model, the average phase of the A1g phonon from Fourier analysis in Fig. 5(b) at the Fermi level is rad. The phase observed here is in good agreement with the previously reported value of rad [15]. It is however worth noting that there is a non-flat variation of the spectral phase across the XUV absorption energy. In the case of a purely displacive excitation, the phase of the A1g phonon should be zero with a small phase imparted by the dampening [15]. The dampening induced phase shift calculated from the phonon dampening and carrier relaxation is found to be less than rad. OḾahony [43] demonstrated in the A7 semimetals that there is a dependence of the phonon phase on lifetime of the electronic excitation, which results in a mixing of a displacive excitation force (flat phase) and an impulsive force that depends on carrier relaxation. The XUV energy dependence of the A1g phonon phase near the Fermi level in Fig. 5(b) thus could be due to such an energy dependent electron-phonon coupling [43]. In the case of a pure displacive excitation, where there is no electronic excitation lifetime dependence, the force exerted on the A1g phonon mode will be constant over the pump pulse and there should be no energy-dependence of the phase. In Te, the measured phase exhibits a small but observable, rad variation. This implies a small, but non-zero effect of carrier relaxation on the A1g phonon driving force. Despite the hot carrier energy thermalization through electron-phonon scattering, which occurs over 0.15(1) ps, this appears to have little impact on the phase. When analyzing the character of the bands excited by the 1.3-2.5 eV pump pulse, this is not surprising. The bandwidth covers the non-bonding lone pair (-3 eV to VBM) and the anti-bonding orbitals (CBM to 3 eV above) crucial to the helical chain structure [44]. All the excited electrons are thus likely to cause a weakening of the intra-chain bonds, as opposed to a strong energy-dependent contribution, which would therefore produce a near pure displacive excitation. This could be tested in detail with theoretical calculations of the energy dependent carrier relaxation as has been done for 2D tellurium [45], or by calculating the band dependence of the force along the A1g phonon coordinate [43].
Although the weak E1TO mode is not distinctively apparent in the time-domain dynamics, due to the small signal strength in the Fourier analysis compared to the dominating A1g mode, we consider the amplitude weighted average phase near the Fermi level (40 to 40.8 eV), where the A is at a maximum and the E1TO mode can be resolved. Unlike the A1g mode attributed to a displacive excitation driven by , the phase at the 2.3-2.5 THz E1TO mode frequency is rad, which is inconsistent with the displacive mechanism. This is not surprising as the chain rotation motion of the E1TO mode is unlikely to be coupled to the lifting of the degeneracy at the Fermi level or population of intra-chain -type anti-bonding orbitals. This is similarly observed in Sb and was attributed to a non-resonant impulsive stimulated Raman scattering [46]. It is also possible to consider excitation of the non-bonding lone pairs of the valence band, which could exert a symmetry breaking force through an asymmetric excited electron distribution, leading to excitation of the E1TO mode [36]. While we found that the coherent phonons are not responsive to changes in laser polarization in the absorption detection geometry with these polycrystalline samples, future studies on samples oriented perpendicular to the c-axis could enhance measurement of the non-totally symmetric modes. It may be interesting to explore whether the E1TO mode may couple to the electronic temperature or influence motion along the A1g coordinate.

### v2

@@ III.3 Temporal Dynamics of the A1 Phonon and Carriers
We now turn to the temporal dynamics of the three SVD components that track the kinetic processes following the excitation. In Fig. 3 the temporal vectors exhibit distinct dynamics that encode the edge shift, broadening, and carriers in detail. Focusing on the exponential rises and decays, the temporal dependence of the edge shift contribution (labeled energy shift) has a sharp initial rise within the first few femtoseconds of the optical excitation followed by a slower onset over a picosecond. The fast rise can be ascribed to a renormalization of the transitions to the conduction band typically driven by changes in the electronic screening of both the valence [35] and 4d core-hole [32, 34]. From the kink in the energy shift trace that occurs briefly after the zero-time delay, this electronic contribution can be estimated to account for nearly one-third of the energy shift. Depopulation of lone pair states near the Fermi level to anti-bonding states in the conduction band [8] also yields rapid destabilization of the ground state potential energy surface resulting in a displacive force along the A1 phonon coordinate that increases the helical chain radius by shifting the equilibrium position (). This increases the symmetry of the Te structure, inducing a displacive excitation of a coherent phonon (DECP) [15] and causes a rise on the order of half the phonon period. The phonon-induced change in local environment is what gives rise to an absorption edge shift due to changes in the 4d binding energy and decrease in the band-gap as discussed above in Section III.2. The picosecond rise is typical of an incoherent growth in the acoustic phonon population that emerges from decay of the optical phonons [36]. The decay of the acoustic phonons is heavily suppressed by the very low lattice thermal conductivity of Te [7] and occurs over the course of milliseconds, as determined from the presence of a large heat background in our TA spectra.
@@
Notably, we observe oscillations in both the energy shift and hot carrier signals. While observation of the A1 phonon in Te is well established, oscillations in the hot carrier signal at the same frequency of the phonon is similar to the back-coupling of the electronic temperature to the lattice that has been observed in semi-metallic Bi [18], Sb [19], and CDW system TaSe2 [20]. The hot carrier signal follows the A1 phonon motion, indicating the coupling arises from movement along the A1 phonon coordinate, corresponding to changes in the Te chain radius. The phase shift implies that when the Te helices are symmetrized (phonon maximum excursion), the carriers have a local temperature minimum. We discuss the potential source of this phase shift and quantify the temporal dynamics in the following section IV.
@@ IV.1 Multi-Temperature Model
To quantify the dynamics of the phonons and carriers the three SVD traces were globally fit weighted by their singular values using (Equation 1) which corresponds to a multi-temperature model. A Gaussian instrument response function, , is retrieved by an external measurement and fixed in the fitting routine. In this model, the hot carriers are an electronic temperature () bath that may transfer energy to the lattice via an electron-phonon scattering rate (). The electronic temperature receives energy via the initially non-thermal carriers following photo-excitation through fast and slow scattering channels consistent with the biexponential rise of the hot carriers (e = electronic and p = phonon). The energy shift is described by the tellurium atom displacement from the equilibrium position along the A1 coordinate (), the acoustic phonon temperature (), and carrier-induced changes in electronic screening (E). The acoustic phonon temperature is fixed to rise at twice the rate of the coherent phonon decay () as the coherent optical phonon has been shown to decay via a Klemen’s like mechanism [36] through an anharmonic coupling of the lattice displacement to degenerate acoustic modes [38, 39, 40]. The rate for the initial rise is described by and relaxation back to the ground state potential energy surface (). The decay of E is shared with in this model, under the approximation they both relate to the carrier population. Finally, the coherent energy exchange between the coherent optical phonon and electronic temperature is described following the energy balance equations coupled to a classical oscillator approach of Giret et al. [17]. This includes the A1 frequency (), coherent phonon phase (), and a frequency chirp of (). The transient broadening in Equation 1 is fit and follows the acoustic phonon temperature and phenomenologically the electronic component of the coherent energy exchange.
@@
At a carrier density of 1.5, the A1 phonon frequency retrieved is = 3.17(1) THz compared to the equilibrium 3.6 THz measured by Raman spectroscopy (see black curves in Fig. 3). The softening of the A1 mode is not surprising as the frequency is known to be dependent upon both carrier density [16, 9], and temperature in the form of anharmonic phonon-phonon coupling [9]. However, when subjected to a few percent excitation of the valence electrons, an increase in temperature is expected to signifigantly blue shift the A1 phonon frequency [9]. The 12 % decrease in frequency of the A1 motion measured with the XUV can then be attributed to a decrease in bond strength as the non-bonding lone pairs at the Fermi level are nearly instantaneously excited into the conduction band anti-bonding orbitals. This is in good agreement with the measured frequency determined from optical reflectivity pump-probe spectroscopy of polycrystalline Te [16, 9]. Similarly, at a carrier density of 2 the A1 phonon frequency is found to be 3.6(3) THz, closer to the equilibrium value (See Appendix F) with no apparent decay. The coherent phonon lifetime of ps and chirp of 0.89(7) ps-2 at 1.5 is typical of high fluence [9], while the coherent phonon phase is determined to be -0.04(1) rad and is discussed further below in Section IV.2.
@@
The rise of the energy shift signal () and decay () of the Te displacement from the equilibrium position () is now considered. The lifetime of the decay of from the excited state potential energy surface as determined by the multi-temperature model is ps and is much slower than the hot carrier decay. This lifetime of the Te displacement in the excited state potential energy surface is nearly identical to the carrier Auger recombination lifetime (31 ps) at the band-edge measured via optical pump-probe spectroscopy in 100 nm thick flakes [14] and the photo-induced degeneracy lifting of the conduction band minimum (H6 band) relaxation time (30 ps) [13]. In the original formulation of the displacive excitation of a coherent phonon model it is unclear whether the displacement is dependent upon the carrier population, , or if only the electronic temperature contributes to the displacement [15]. It is notable that in the multi-temperature model follows a timescale comparable to Auger recombination of fully relaxed carriers, such that . This could potentially explain why the displacement of the Te equilibrium position is much larger than in A7 semimetals [19, 18] where it is proposed to follow the electronic temperature. Additionally, when measured at a carrier density of , has no measurable decay out to the maximum time delays tested ( 2 ps). This is consistent with lengthening of the carrier lifetime as Auger recombination is reduced at low carrier densities, and multi-phonon mediated recombination is only expected to occur over 200 ps [14]. Moreover, there is a fast rise () to the energy shift trace, which is instantaneous with the sub 5 fs pump excitation (see Appendix G). This rise, while too fast to be explained by the phonon motion, likely originates in carrier-induced electronic structure changes such as changes in electron localization [34] or 4d core-hole screening [32].
@@
The dynamics of the hot carriers are examined in Fig. 3. Upon initial excitation, the non-thermal carriers far from-equilibrium will quickly redistribute their energy through various scattering channels as they thermalize. The biexponential rise of the multi-temperature model for the hot carriers includes a fast term that closely follows the 6.4(5) fs instrument response, and a slower 196(8) fs rise. The latter is extracted by deconvoluting the rise dynamics from the cross-correlation, while the fast term is not readily resolved. Both timescales are much faster than previously measured inter-valley scattering that necessitates a population of high momentum acoustic phonons [41, 13]. Thus the primary source is likely intra-valley scattering towards the Fermi level in the form of energy-relaxation that is typically on the order of 100 fs [42, 43]. It is however possible that both carrier-carrier and carrier-optical phonon scattering are the dominant mechanisms for thermalization [44, 27]. While carrier-carrier scattering can only redistribute energy within the electronic system, the maximum energy exchange is on the order of the 1.55 eV photon energy. This will lead to much faster thermalization than through optical phonon emission that is constrained to the tens of meV transfer per scattering event. The fast rise () is thus attributable to carrier-carrier scattering, while the slower thermalization channel can be attributed to electron-phonon coupling, which is similar to the 0.3 ps intra sub-band cooling time measured with photoelectron spectroscopy [11]. The discrepancy could be due to the fact that in the analysis the hot carrier signal included both holes and electrons and that there are differences in scattering rates for holes versus electrons. The hot carrier decay () has a relaxation time of 1.59(3) ps that is similar to the timescale (1.3 ps [41], 1.7 ps [11]) for the transfer of high-energy carriers to the lowest energy sub-bands of the H-valley. In this process, for the hot carrier decay, the electronic temperature is reduced as optical-phonon emission through electron-phonon coupling that transfers energy to the lattice. The hot carriers determined from the SVD do not include the cooled lowest energy carriers that share the long lifetime of the Te equilibrium position displacement. Similarly, the non-thermal carriers are shown in the residual TA spectra alongside the cooled carriers (See Fig. 10 of Appendix E), where there are residual changes in absorption near the Fermi-level that rapidly decay on a timescale of a hundred femtoseconds. This underlies a clear limitation of the analysis above. The SVD is not sensitive to early timescale temporal variations in the spectral profile (on the order of a mOD), such as the non-thermal carriers and their deviation from a Fermi-Dirac distribution in the first tens of femtoseconds. Thus, similar to other XUV TA studies, the thermalization timescales are retrieved by analyzing the appearance of a maximum in the hot carrier population, as non-thermal carriers may produce only small spectral deviations [45].
IV.2 Coupled Carrier and A1 Phonon Dynamics
@@
We now consider the coherent modulations of that are observed. Following the absorption of an ultrashort optical laser pulse by Te, there is a rapid increase in the total energy of the electronic degrees of freedom. This abrupt increase in the carrier population () and electronic temperature () contributes to excite the A1 phonon mode and is well explained by the DECP model [15] and additional theory by O’Mahony et al. [46]. While Te is a semiconductor in the ground state, the displacement along the A1 phonon coordinate leads to a symmetric Dirac metal through a light-induced phase transition [8]. This phase transition can similarly be achieved by depopulating the doubly degenerate VBM lone-pair orbitals with shear strain to the CBM [3]. The increase in metallic character when Te becomes more symmetric by increasing the helical chain radius should lead to an increase in the electronic heat capacity [47] as the DOS is increased at the Fermi level [31]. If we consider the temporal evolution of the hot carriers in Fig. 3, the phase offset is close to with respect to the phonon excursion. When the phonon displacement is maximum and Te most metallic, the electronic heat capacity is maximized, leading to a lower temperature of the hot carriers. This similarly follows the thermodynamic model of Giret et al. [17] that describes a coupling between the electronic temperature and lattice motion in Bi as an isentropic conservation of entropy. The same model has been applied to interpret XUV pump-probe results in Sb [19] and Bi [18]. As shown by Ning et al. [8], A1 coherent phonon motion leads to vast modifications of the band structure in the vicinity of the Fermi level, and at a maximum distortion, metallization of Te. Given this, the dependence of the electronic heat capacity on could exhibit large variation due to how it strongly couples to the band structure, particularly if there is a large excursion of the A1 phonon mode (given as a radius). This is not surprising as there are numerous band crossings at the Fermi level that emerge upon increasingly symmetrizing Te [8, 31]. While an indirect band gap is closed at the A-valley for a 14 pm displacement [31], complete symmetrizing of Te at 26 pm additionally closes the gap at the H-valley [8]. Although our level of analysis precludes determination of a band gap closing, or the extent of symmetrization, the oscillations of the hot carrier are indicative of a strong influence of the A1 phonon on the electronic heat capacity, and thus the transiently changing number of available states in proximity to the Fermi level. Calculations of the electronic heat capacity changes that occur from the light-induced excitation of the A1 phonon mode may prove useful in the future in understanding these observations, and to determine which valleys may contribute the most for the carriers excited.
@@
Finally, we consider a Fourier analysis of the phonon-induced oscillations shown in Fig. 4. This is accomplished by subtracting the exponential contributions to the TA captured by the SVD and adding back the residual TA of Fig. 10. Further, a decomposition of the spin-orbit split 4d core-levels is done following the methodology of Zuerch et al. [30] to eliminate spectral congestion between the and to conduction band transitions. Fig. 4(a), the power spectrum () with only the phonon-induced oscillations shows a maximum intensity near the Fermi level of 40.4 eV, while above 42 eV, weaker features are associated with the appearance of phonon-induced features from the large lattice distortion found in Fig. 2(b).
@@
In Fig. 4(b), the spectral phase is shown for the A1 mode in vicinity of the Fermi level. In agreement with the phase of rad retrieved from the multi-temperature model, the average phase of the A1 phonon from Fourier analysis in Fig. 4(b) at the Fermi level is rad. The phase observed here is in good agreement with the previously reported value of rad [15]. It is however worth noting that there is a non-flat variation of the spectral phase across the XUV absorption energy. In the case of a purely displacive excitation, the phase of the A1 phonon should be zero with a small phase imparted by the dampening [15]. The dampening induced phase shift calculated from the phonon dampening and carrier relaxation is found to be less than rad. OḾahony [48] demonstrated in the A7 semimetals that there is a dependence of the phonon phase on lifetime of the electronic excitation, which results in a mixing of a displacive excitation force (flat phase) and an impulsive force that depends on carrier relaxation. The XUV energy dependence of the A1 phonon phase near the Fermi level in Fig. 4(b) thus could be due to such an energy dependent electron-phonon coupling [48]. In the case of a pure displacive excitation, where there is no electronic excitation lifetime dependence, the force exerted on the A1 phonon mode will be constant over the pump pulse and there should be no energy-dependence of the phase. In Te, the measured phase exhibits a small but observable, rad variation. This implies a small, but non-zero effect of carrier relaxation on the A1 phonon driving force. Despite the hot carrier energy thermalization through electron-phonon scattering, which occurs over 196(8) fs, this appears to have little impact on the phase. When analyzing the character of the bands excited by the 1.3-2.5 eV pump pulse, this is not surprising. The bandwidth covers the non-bonding lone pair (-3 eV to VBM) and the anti-bonding orbitals (CBM to 3 eV above) crucial to the helical chain structure [49]. All the excited electrons are thus likely to cause a weakening of the intra-chain bonds, as opposed to a strong energy-dependent contribution, which would therefore produce a near pure displacive excitation. This could be tested in detail with theoretical calculations of the energy dependent carrier relaxation as has been done for 2D tellurium [50], or by calculating the band dependence of the force along the A1 phonon coordinate [48].
While we found that the A1 coherent phonon and hot carriers are not responsive to changes in laser polarization in the absorption detection geometry with these polycrystalline samples, future studies on samples oriented perpendicular to the c-axis could enable measurement of the non-totally symmetric modes through the photo-Dember effect [51]. It may be interesting to explore whether the E1TO mode or other phonon modes may couple to the electronic temperature or influence motion along the A1 coordinate.